OpenStudy (anonymous):

What is the exact length of the unknown side? A right triangle appears. One leg is labeled seven. One leg is labeled with a question mark. The hypotenuse is labeled nine. the square root of twenty the square root of thirty-two the square root of sixty-three the square root of one hundred thirty

OpenStudy (ajprincess):

 Use pythagoras theorem. \(a^2+b^2=c^2\) \(7^2+a^2=9^2\) \(49+a^2=81\)